The Ksp for silver(I) phosphate is 1.8 x 10–18. Calculate the molar solubility of silver(I) phosphate.

Group of answer choices

1.8 x 10–1 M

1.6 x 10–5 M

2.1 x 10–5 M

7.2 x 10–1 M

3.7 x 10–5 M
Answer #1

Solution:-

Ag3PO4 (s)  \rightleftharpoons 3 Ag+ (aq) + PO43- (aq)

Let, the solubility of Ag​​​​​​​3​​​​​​​PO4 be "s" M.

[Ag​​​​​​​+] = 3s

[PO43-] = s

Thus, the solubility product,

Ksp = [Ag​​​​​​​+]3 [PO43-]

(1.8 × 10-18) = (3s)3 (s) = 27s4

or, s = (1.8 × 10-18 / 27)1/4 = 1.6 × 10-5 M

Hence, molar solubility of silver(I) phosphate is = 1.6 × 10-5 M

------------------------------------------------------------------------------